OpenStudy (anonymous):

In physics what is a field? A) the immediate space around a charge, mass or magnet b) a region in which something experiences a force C) the space in which objects are able to move D) The region between two charges, masses or magnets

OpenStudy (anonymous):

I think its B but im not 100% positive

OpenStudy (anonymous):

I would say C.

OpenStudy (anonymous):

the larger the object the more field lines? is that true?

OpenStudy (anonymous):

A field is associated with the plotting of spacetime. C is the only answer that is general enough to cover this fact. Generally speaking, yes, the larger the object the "more" field there is.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

... Proof, for those who disagree. http://en.wikipedia.org/wiki/Field_%28physics%29

OpenStudy (anonymous):

Define the concept of a gravitational field?

OpenStudy (anonymous):

It's a field that is caused due to the bump in spacetime caused by mass. Sort of hard to imagine in 3D.

OpenStudy (anonymous):

That is by no means a definition. It's just to give you an idea.

OpenStudy (anonymous):

A field where the acceleration of gravity acts as a force on all bodies inside. Alternatively, a field that is a geodesic, in terms of general relativity.

OpenStudy (anonymous):

Doesn't he mean the origin of gravity?

OpenStudy (anonymous):

"The concept of a gravitational field". The field itself, not the mass. But you covered for that anyways, so I'm not worried. ;D

OpenStudy (anonymous):

What would be the gravitational field at the midpoint between, and due to two identical planets? explain

OpenStudy (anonymous):

Be a zero net force. Nearly impossible in reality. It would just stay in position.

OpenStudy (anonymous):

Thanks guys! I have a few more questions :)

OpenStudy (anonymous):

An object of mass 12kg weighs 156 N on planet X. what is the magnitude of the acceleration due to gravity on planet x

OpenStudy (anonymous):

F = ma a = F/m 156/12 = something

OpenStudy (anonymous):

What LordHades said. I'm just piggybacking his medals.

OpenStudy (anonymous):

how might this planet differ from Earth?

OpenStudy (anonymous):

Well, its acceleration due to gravity is much higher, so it's probably of a greater mass.

OpenStudy (anonymous):

And you;ll weigh heavier there.

OpenStudy (anonymous):

12*9.81= 117.6N

OpenStudy (anonymous):

Note that I'm talking about weight not mass.

OpenStudy (anonymous):

that would be on earth, so you are lighter on earth

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

The Earth has a mass of 6x10^24kg and a diameter of 12,765 km. A 500kg satellite experiences a gravitational acceleration of 0.4 ms^-2 in its orbit. What is the altitude of the satellite?

OpenStudy (anonymous):

M = 6 x 10^24 re = 6382.5 m = 500 ac = 0.4 ac = v^2/r r = v^2 / ac -------->1 GMm/r^2 = mv^2/r ------------>2 From 1 and 2, Solve for 'v' and then get 'r' Should work I think, tell me if I messed up somewhere.

OpenStudy (anonymous):

List of possible answers a) 2.53 x10^7 m b) 3.16 x 10^7 m c) 2.80 x 10^8 m d) 2.86 x 10^8m

OpenStudy (anonymous):

From here on it's just calculations and algebra, nothing difficult.

OpenStudy (anonymous):

At what altitude would a 1000kg satellite be to experience the same gravitational acceleration? Explain

OpenStudy (anonymous):

You could use the exact same formulae, just need to rearrange the variables.

OpenStudy (anonymous):

I dont have a calculator with me :(

OpenStudy (anonymous):

Doubling the mass, yet keeping the acceleration the same, is multiplying both sides of the equation by 2.

OpenStudy (anonymous):

The force vector of the satellite, the one we're concerned with, doubles, right? Yeah. The mass of the satellite also doubles.

OpenStudy (unklerhaukus):

do you know about the Lagrange points in a feild?

OpenStudy (anonymous):

no... what are they?

OpenStudy (unklerhaukus):

the stable points, for example in the sun earth system there are 5 stable points (places where gravitational attraction is balanced)